OpenStudy (anonymous):

How does one derive sinc(0)=1 from lim(sinx/x)=1? Would not sinc(0)->1 be a lot more accurate without explicitly defining that sinc(0)=1?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sin(x)/x = 1\] However, the function \[\sin(x)/x\] is not defined in x = 0 (denominator cnanot be 0) To make \[sinc(x)\] defined at any point, we say\[sinc(x)\] is 1.